# Gated Integration of Low-Rank Adaptation for Continual Learning of Language Models

### v1

@@ C.4.3 Varying the Layers of Gating Module
C.5 Varying Ranks in LoRA Branches
@@
C.6 Adopting Other Update Strategies for the New LoRA Branch
Our GainLoRA does not impose specific update strategies for the new LoRA branches. In this work, we adopt the same update strategies as the existing two methods, O-LoRA wang2023orthogonal and InfLoRA liang2024inflora . Related methods, such as IncLoRA hu2021lora and C-LoRA smithcontinual , also adopt the expandable LoRA architecture illustrated in Figure 1 (a) and fix all integration coefficients $\{a_{i}\}_{i=1}^{T}$ to 1. Our method GainLoRA can also adopt their update strategies for the new LoRA branch. Table 13 presents the results, demonstrating that GainLoRA further improves the performance of these two methods.
@@ C.6 Adopting Other Update Strategies for the New LoRA Branch
IncLoRA 12.33$\pm$0.56 41.93$\pm$0.17 16.65$\pm$0.91 36.56$\pm$1.30
GainLoRA (IncLoRA) 47.82$\pm$0.08 3.73$\pm$0.25 45.42$\pm$1.19 5.83$\pm$1.53
C-LoRA 22.69$\pm$0.01 24.25$\pm$0.90 32.81$\pm$0.64 11.60$\pm$0.23
GainLoRA (C-LoRA) 49.24$\pm$0.21 2.94$\pm$0.41 46.23$\pm$0.61 6.05$\pm$0.51
C.7 Extending to the Rehearsal Setting
In this work, we focus on the non-rehearsal setting, where no real or synthetic samples from old tasks are available during the learning of a new task. In this section, we demonstrate that our method, GainLoRA, can also be extended to the rehearsal setting. Specifically, in the rehearsal setting, a set of samples $\mathcal{N}_{t}$ containing real or synthetic samples from the previous $t-1$ tasks is available while the model learns the $t$-th new task. In this case, the constraints introduced in Section 3.2.1 are no longer necessary, and we can optimize the following loss function:
$\frac{1}{|\mathcal{D}_{t}|}\sum_{(\bm{x}_{t},\bm{y}_{t})\in\mathcal{D}_{t}}\sum_{j=1}^{|\bm{y}_{t}|}\log\left[P(y_{t,j}|\bm{x}_{t},y_{t,1},...,y_{t,j-1})\right]+\sum_{(\bm{x},\bm{y})\in\mathcal{N}_{t}}\log g_{t}(\bm{x}),$ (22)
The second term in (22) minimizes the contribution from the new LoRA branch on old tasks.
We compare GainLoRA with SAPT-LoRA zhao2024sapt . For a fair comparison, we use the same rehearsal dataset as SAPT-LoRA, generated using a trained generative model. As shown in Table 14, GainLoRA achieves comparable performance to SAPT-LoRA in the rehearsal setting. Note that SAPT-LoRA is specifically designed for the rehearsal setting and is not applicable to the non-rehearsal setting, which is considered in this work.
Table 14: The overall results on Order 1 in the rehearsal-setting.
T5-Large Llama-2-7B
[…]
SAPT-LoRA zhao2024sapt 51.38$\pm$0.12 0.74$\pm$0.18 55.88$\pm$0.25 0.74$\pm$0.27
GainLoRA (InfLoRA) + Replay 51.62$\pm$0.56 0.08$\pm$0.10 55.93$\pm$0.69 0.95$\pm$0.39
C.8 Performance on the TRACE Benchmark
To further demonstrate our method’s effectiveness, we follow existing CL methods for LMs he2024seekr ; wang2023trace and conduct experiments on the TRACE dataset wang2023trace with Llama-2-7B-Chat. The dataset comprises a diverse set of challenging instruction-tuned tasks, spanning multilingual comprehension, domain-specific knowledge, arithmetic reasoning, and coding. An overview of the tasks in TRACE is provided in Table 15.
@@ C.8 Performance on the TRACE Benchmark
TRACE Benchmark C-STANCE $\rightarrow$ FOMC $\rightarrow$ MeetingBank $\rightarrow$ Py150 $\rightarrow$ ScienceQA $\rightarrow$ NumGLUE-cm $\rightarrow$ NUMGLUE-ds $\rightarrow$ 20Minuten
@@
We also follow existing work wang2023trace to explicitly evaluate the preservation of general abilities, such as instruction-following, after continual learning on the TRACE benchmark using different methods. Table 17 indicates that continual learning with different methods often leads to a degradation of general abilities. However, GainLoRA demonstrates a stronger ability to mitigate forgetting compared to other LoRA-based methods, including O-LoRA and InfLoRA.
C.9 Compared with More CL Methods in CV
Following many existing continual learning methods in NLP zhao2024sapt ; wang2023orthogonal ; he2024seekr , this paper focuses on models based on next-token prediction, such as T5 raffel2020exploring and LLaMA touvron2023llama2 , which lack the [CLS] token used in ViT. Although many continual learning methods based on pre-trained models in computer vision wang2022dualprompt ; wang2022learning ; wang2022s ; wang2023hierarchical ; wang2024hide ; smith2023coda cannot be directly applied to our setting, we adapt several of them to the T5 architecture to ensure a comprehensive comparison. Specifically, we select three representative methods from the vision domain, including L2P wang2022learning , DualPrompt wang2022dualprompt , and CODA-Prompt smith2023coda . These methods rely on the [CLS] token in ViT to retrieve prompts from a prompt pool for training and inference. Since T5 does not provide a [CLS] token, we instead use the mean representation of each input sequence as a query. We carefully tuned the learning rate, epochs, and their unique hyperparameters on the validation set. Table 18 presents the performance of these adapted methods on order 1 with T5. As shown, they perform significantly worse than our GainLoRA and exhibit noticeable forgetting.
@@ C.9 Compared with More CL Methods in CV
L2P wang2022learning 14.59 24.70
DualPrompt wang2022dualprompt 11.77 12.95
CODA-Prompt smith2023coda 10.80 16.15

### v2

@@ C.3.3 Varying the Layers of Gating Module
C.4 Varying Ranks in LoRA Branches
@@
C.5 Adopting Other Update Strategies for the New LoRA Branch
Our GainLoRA does not impose specific update strategies for the new LoRA branches. In this work, we adopt the same update strategies as the existing two methods, O-LoRA [61] and InfLoRA [32]. Related methods, such as IncLoRA [21] and C-LoRA [50], also adopt the expandable LoRA architecture illustrated in Figure 1 (a) and fix all integration coefficients $\{a_{i}\}_{i=1}^{T}$ to 1. Our method GainLoRA can also adopt their update strategies for the new LoRA branch. Table 10 presents the results, demonstrating that GainLoRA further improves the performance of these two methods.
@@ C.5 Adopting Other Update Strategies for the New LoRA Branch
IncLoRA 12.33 41.93 16.65 36.56
GainLoRA (IncLoRA) 47.82 3.73 45.42 5.83
C-LoRA 22.69 24.25 32.81 11.60
GainLoRA (C-LoRA) 49.24 2.94 46.23 6.05
C.6 Performance on the TRACE Benchmark
To further demonstrate our method’s effectiveness, we follow existing CL methods for LLMs [18, 62] and conduct experiments on the TRACE dataset [62] with Llama-2-7B-Chat. The dataset comprises a diverse set of challenging instruction-tuned tasks, spanning multilingual comprehension, domain-specific knowledge, arithmetic reasoning, and coding. An overview of the tasks in TRACE is provided in Table 11.
@@ C.6 Performance on the TRACE Benchmark
TRACE Benchmark
C-STANCE $\rightarrow$ FOMC $\rightarrow$ MeetingBank $\rightarrow$ Py150 $\rightarrow$
ScienceQA $\rightarrow$ NumGLUE-cm $\rightarrow$ NUMGLUE-ds $\rightarrow$ 20Minuten
@@
We also follow existing work [62] to explicitly evaluate the preservation of general abilities, such as instruction-following, after continual learning on the TRACE benchmark using different methods. Table 13 indicates that continual learning with different methods often leads to a degradation of general abilities. However, GainLoRA demonstrates a stronger ability to mitigate forgetting compared to other LoRA-based methods, including O-LoRA and InfLoRA.
C.7 Compared with More CL Methods in CV
Following many existing continual learning methods in NLP [75, 61, 18], this paper focuses on models based on next-token prediction, such as T5 [43] and LLaMA [54], which lack the [CLS] token used in ViT. Although many continual learning methods based on pre-trained models in computer vision [66, 67, 63, 57, 58, 51] cannot be directly applied to our setting, we adapt several of them to the T5 architecture to ensure a comprehensive comparison. Specifically, We implement these methods by injecting prompts into both the keys and values in T5, and introduce zero-padding in the position bias tensor to ensure shape compatibility. Note that we do not add positional information to the prompts, which is consistent with DualPrompt and CODA-Prompt in ViT. The results on order 1 with T5 are reported in Table 14, showing that these methods perform significantly worse than our GainLoRA and exhibit noticeable forgetting.
@@ C.7 Compared with More CL Methods in CV
L2P [67] 15.23 11.34
DualPrompt [66] 17.40 10.63
CODA-Prompt [51] 19.28 14.62